# On von Neumann regularity of ample groupoid algebras

### v2

@@ 2.4. Regular rings
A ring $R$ is von Neumann regular, or regular for short, if for all $a\in R$, there exists $b\in R$ with $aba=a$. We do note require rings to be unital. The following facts are quite standard [19].
@@
    The direct product of regular rings is regular.
@@
    If $R$ is a ring and $I$ is an ideal of $R$, then $R$ is regular if and only if $I$ and $R/I$ are regular.
@@
    A ring $R$ is regular if and only if it has local units and each unitary $R$-module is flat.
Recall that a semigroup is locally finite if all its finitely generated subsemigroups are finite, or, equivalently, it is a direct limit of finite semigroups. Note that any subsemigroup of a locally finite semigroup is locally finite. The following is a special case of a classical theorem of Birkhoff in universal algebra, cf. [8, Theorem 10.16]. We include a proof for the sake of completeness.
@@
Let $G$ be a group and $R$ a ring. Then the group ring $RG$ is regular if and only if $R$ is regular, $G$ is locally finite and the order of each element of $G$ is a unit in $R$.
Attempts have been made to extend this to inverse semigroups. The best result to date is in Okniński [27, Chapter 15].
Theorem 2.4.
Let $K$ be a field and $S$ an inverse semigroup.
  1. (1)
    If $K$ has characteristic 00, then $KS$ is regular if and only if $S$ is locally finite.
  2. (2)
    If $K$ has characteristic $p>0$, $S$ is locally finite and $p$ does not divide the order of any element of any maximal subgroup of $S$, then $KS$ is regular.
  3. (3)
    If $KS$ is regular and $K$ has characteristic $p>0$, then $S$ is periodic and each maximal subgroup of $S$ is locally finite with no element of order divisible by $p$.
It is not known in positive characteristic if $KS$ regular implies that $S$ is locally finite for inverse semigroups in general, although this is known for Clifford semigroups, cf. [28, Chapter 15, Proposition 11].
Since we wish to prove the difficult implications of Theorem 2.4 as consequences of our main results, we prove the straightforward implications of Theorem 2.4, which we shall need in our arguments.
@@
Let $K$ be a field and $S$ a locally finite inverse semigroup such that the characteristic of $K$ does not divide the order of any element of any maximal subgroup of $S$. Then $KS$ is regular.
@@
Since the class of regular rings is closed under direct limits by Proposition 2.1, we may assume that $S$ is a finite inverse semigroup and $K$ is a field whose characteristic does not divide the order of any maximal subgroup of $S$. But then $KS$ is semisimple (in fact, isomorphic to a direct product of matrix algebras over group algebras of maximal subgroups) and hence regular, cf. [32, Corollary 9.4]. ∎
3. von Neumann regularity of Steinberg algebras
The following result is proved in [4]. We provide a proof for the reader’s convenience.
Theorem 3.1.
Let $K$ be a field and $\mathscr{G}$ an ample groupoid. If $K\mathscr{G}$ is regular, then each isotropy group of $\mathscr{G}$ is locally finite and the characteristic of $K$ divides the order of no isotropy group element.
Proof.
By Connell’s theorem, it is enough to show that $K\mathscr{G}_{x}^{x}$ is regular for each $x\in\mathscr{G}^{(0)}$. Let $a\in K\mathscr{G}_{x}^{x}$, which we view as a finitely supported function $\mathscr{G}_{x}^{x}\to K$. Choose a compact open bisection $U_{\gamma}$ containing $\gamma$ for each $\gamma\in\mathscr{G}_{x}^{x}$ with $a(\gamma)\neq 0$. Consider $f=\sum_{\gamma\in\mathscr{G}^{x}_{x}}a(\gamma)U_{\gamma}$. By construction, $f|_{\mathscr{G}_{x}^{x}}=a$. Since $K\mathscr{G}$ is regular, we can find $g\in K\mathscr{G}$ with $fgf=f$. Notice that if $\alpha\in\mathscr{G}_{x}^{x}$, then
$f(\alpha)=\sum_{\beta\gamma\nu=\alpha}f(\beta)g(\gamma)f(\nu).$ (3.1)
But if $\alpha=\beta\gamma\nu$, then by construction of $f$ and the definition of a bisection if $f(\beta)\neq 0\neq f(\nu)$, we must have $\beta,\nu\in\mathscr{G}^{x}_{x}$ and hence $\gamma\in\mathscr{G}^{x}_{x}$. Therefore, (3.1) becomes
$a(\alpha)=f(\alpha)=\sum_{\beta,\gamma,\nu\in\mathscr{G}_{x}^{x},\beta\gamma\nu=\alpha}a(\beta)g(\gamma)a(\nu).$
Thus if we put $b=g|_{\mathscr{G}_{x}^{x}}$, then $b$ has finite support and $aba=a$. ∎
We aim to improve on this, particularly in characteristic zero, where we can use some analysis. We shall also see that the necessary condition in Theorem 3.1 is by no means sufficient.
Let us say that an ample groupoid $\mathscr{G}$ is uniformly bounded if there is an integer $M\geq 1$ with $|\mathop{\boldsymbol{r}}^{-1}(x)|\leq M$ for all $x\in\mathscr{G}^{(0)}$. Note that any open or closed subgroupoid of a uniformly bounded ample groupoid is uniformly bounded. Let us say that $\mathscr{G}$ is approximately uniformly bounded if it is a directed union of uniformly bounded open subgroupoids. An open or closed subgroupoid of an approximately uniformly bounded ample groupoid is again approximately uniformly bounded. Let us say that $\mathscr{G}$ is approximately quasi-compact if $\mathscr{G}$ is a directed union of quasi-compact open subgroupoids. It is not difficult to verify that a principal second countable Hausdorff ample groupoid is approximately quasi-compact if and only if it is an approximately finite groupoid in the sense of [26, Section 11.5] (i.e., an approximately finite étale equivalence relation in the sense of [18]). Note that if $\mathscr{G}$ is a quasi-compact ample groupoid, then $\mathscr{G}^{(0)}$ is compact being the image of $\mathscr{G}$ under $\mathop{\boldsymbol{r}}$.
@@
Let $C$ be a quasi-compact subspace of $\mathscr{G}$ containing $\mathscr{H}$. We may cover $C$ by finitely many compact open bisections $U_{1}\cup\cdots\cup U_{n}$ of $\mathscr{G}$. Then we have $|\mathop{\boldsymbol{r}}^{-1}(x)\cap\mathscr{H}|\leq n$ for all $x\in\mathscr{H}^{(0)}$ as $|\mathop{\boldsymbol{r}}^{-1}(x)\cap U_{i}|\leq 1$ for $i=1,\ldots,n$ and $\mathscr{H}\subseteq U_{1}\cup\cdots\cup U_{n}$. ∎
@@
Suppose that (2) holds. If $\mathscr{G}$ is the directed union of uniformly bounded open subgroupoids $\mathscr{G}_{\alpha}$ with $\alpha\in D$, then $\Gamma_{c}(\mathscr{G})$ is the directed union of the $\Gamma_{c}(\mathscr{G}_{\alpha})$ since any compact set belongs to some $\mathscr{G}_{\alpha}$. Hence any finitely generated subsemigroup of $\Gamma_{c}(\mathscr{G})$ is contained in some $\Gamma_{c}(\mathscr{G}_{\alpha})$. Thus we may assume without loss of generality that $\mathscr{G}$ is uniformly bounded. Suppose that $|\mathop{\boldsymbol{r}}(x)|\leq n$ for all $x\in\mathscr{G}^{(0)}$. Then, for each $x\in\mathscr{G}^{(0)}$, we have that $|\mathscr{G}^{x}_{x}|\cdot|\mathcal{O}_{x}|=|\mathop{\boldsymbol{r}}^{-1}(x)|\leq n$. It follows that $|\mathscr{G}|_{\mathcal{O}_{x}}|=|\mathcal{O}_{x}|^{2}\cdot|\mathscr{G}_{x}^{x}|\leq n^{2}$, and hence $\mathscr{G}|_{\mathcal{O}_{x}}$ is a finite closed subgroupoid and $\Gamma_{c}(\mathscr{G}|_{\mathcal{O}_{x}})$ is a finite inverse semigroup of cardinality at most $2^{n^{2}}$.
@@
To prove our main result, let us first recall the construction of the reduced $C^{*}$-algebra of an ample groupoid as described in [11]. Let $\mathscr{G}$ be an ample groupoid. Then $\mathbb{C}\mathscr{G}$ is a $\ast$-algebra with involution $f^{\ast}(\gamma)=\overline{f(\gamma^{-1})}$. For each $x\in\mathscr{G}^{(0)}$, we consider a $\ast$-representation $\pi_{x}\colon\mathbb{C}\mathscr{G}\to\mathcal{B}(\ell_{2}(\mathscr{G}_{x}))$ given on the orthonormal basis $\delta_{\gamma}$, $\gamma\in\mathscr{G}_{x}$, by
$\pi_{x}(f)\delta_{\gamma}=\sum_{\alpha\in\mathop{\boldsymbol{s}}^{-1}(\mathop{\boldsymbol{r}}(\gamma))}f(\alpha)\delta_{\alpha\gamma}.$
Notice that if $U$ is a compact open bisection, then $\|\pi_{x}(1_{U})\|\leq 1$. Indeed, $U$ induces a partial permutation $\psi_{U}$ of $\mathscr{G}_{x}$ by observing that if $\gamma\in\mathscr{G}_{x}$, then either $U\{\gamma\}$ is empty, in which case $\psi_{U}(\gamma)$ is undefined, or $U\{\gamma\}=\{\alpha\gamma\}$ where $\alpha$ is the unique element of $U$ with $\mathop{\boldsymbol{s}}(\alpha)=\mathop{\boldsymbol{r}}(\gamma)$, in which case $\psi_{U}(\gamma)=\alpha\gamma$. Therefore if $a=\sum_{\gamma\in\mathscr{G}_{x}}c_{\gamma}\delta_{\gamma}\in\ell_{2}(\mathscr{G}_{x})$, then writing $\mathrm{dom}(\psi_{U})$ for the domain of $\psi_{U}$, we have that
$\pi_{x}(1_{U})a=\sum_{\gamma\in\mathscr{G}_{x}}\sum_{\alpha\in\mathop{\boldsymbol{r}}^{-1}(\gamma)}1_{U}(\alpha)c_{\gamma}\delta_{\alpha\gamma}=\sum_{\gamma\in\mathrm{dom}(\psi_{U})}c_{\gamma}\delta_{\psi_{U}(\gamma)}$
and so $\|\pi_{x}(1_{U})a\|\leq\|a\|$. Thus $\|\pi_{x}(1_{U})\|\leq 1$. It follows that if $f=\sum_{i=1}^{n}c_{i}1_{U_{i}}\in\mathbb{C}\mathscr{G}$ with the $U_{i}$ compact open bisections, then $\|\pi_{x}(f)\|\leq\sum_{i=1}^{n}|c_{i}|$, and so $\|\pi_{x}(f)\|\leq|f|$ where ||$\cdot$|| ⋅ | is the norm defined in [33] by
$|f|=\inf\left\{\sum_{i=1}^{n}|d_{i}|:f=\sum_{i=1}^{n}d_{i}1_{U_{i}},U_{i}\in\Gamma_{c}(\mathscr{G})\right\}.$
(Note that this latter norm also bounds the $I$-norm.) The reduced $C^{*}$-norm on $\mathbb{C}\mathscr{G}$ is defined by $\|f\|_{r}=\sup_{x\in\mathscr{G}^{(0)}}\|\pi_{x}(f)\|$. The reduced $C^{*}$-algebra $C^{*}_{r}(\mathscr{G})$ of $\mathscr{G}$ is then the completion of $\mathbb{C}\mathscr{G}$ with respect to $\|\cdot\|_{r}$; see [11] for details. The fact that the reduced norm is truly a norm, and hence $\mathbb{C}\mathscr{G}$ embeds in $C^{*}_{r}(\mathscr{G})$, follows from the following estimate that we shall need in any event.
@@
If $f\in\mathbb{C}\mathscr{G}$, then $\|f\|_{\infty}\leq\|f\|_{r}\leq|f|$.
@@
We already observed that $\|f\|_{r}\leq|f|$. If $f\in\mathbb{C}\mathscr{G}$ and $\gamma\in\mathscr{G}$, then for $x=\mathop{\boldsymbol{s}}(\gamma)$, we have $\langle\pi_{x}(f)\delta_{x},\delta_{\gamma}\rangle=f(\gamma)$, and so by the Cauchy-Schwarz inequality $|f(\gamma)|\leq\|\pi_{x}(f)\|\leq\|f\|_{r}$. ∎
We now aim to characterize ample groupoids with regular algebras over fields of characteristic zero, generalizing Theorem 2.4. Our proof is inspired by Okniński [28, Chapter 13, Lemma 2], who traces his approach to an idea of Formanek. We shall need the well-known fact that any countable field $L$ of characteristic 00 can be embedded into $\mathbb{C}$. Indeed, choose a transcendence basis $S$ for $L/\mathbb{Q}$. Then $L/\mathbb{Q}(S)$ is algebraic and $\mathbb{Q}(S)/\mathbb{Q}$ is purely transcendental. Since $L$ is countable, $S$ is also countable. On the other hand, $\mathbb{C}$ is uncountable and hence has an uncountable transcendence basis $T$ over $\mathbb{Q}$. Thus we can find an injective map $S\to T$ which extends to a field embedding $\mathbb{Q}(S)\hookrightarrow\mathbb{C}$. Since $L/\mathbb{Q}(S)$ is algebraic and $\mathbb{C}$ is algebraically closed, we can then extend this embedding to $L$.
Theorem 3.7.
Let $K$ be a field of characteristic zero and $\mathscr{G}$ an ample groupoid. Then the following are equivalent.
  1. (1)
    $K\mathscr{G}$ is regular.
  2. (2)
    $\mathscr{G}$ is approximately uniformly bounded.
  3. (3)
    $\mathscr{G}$ is approximately quasi-compact.
  4. (4)
    $\Gamma_{c}(\mathscr{G})$ is locally finite.
  5. (5)
@@
We already know that (2)–(5) are equivalent by Proposition 3.4. Let $S=\Gamma_{c}(\mathscr{G})$ be the inverse semigroup of compact open bisections of $\mathscr{G}$. Note that (4) implies that $KS$ is regular by Proposition 2.5, and hence $K\mathscr{G}$ is regular by Proposition 2.1, as it is quotient of $KS$. Thus (4) implies (1). It remains to show that (1) implies (2).
Note that if $U$ is any compact open subset of $\mathscr{G}^{(0)}$, then $U\mathscr{G}U=\mathscr{G}|_{U}$ is an open subgroupoid with compact unit space $U$ and $K\mathscr{G}|_{U}=1_{U}K\mathscr{G}1_{U}$, and hence is regular by Proposition 2.1. Since $\mathscr{G}$ is the directed union of the $\mathscr{G}|_{U}$ as $U$ runs over all compact open subsets of $\mathscr{G}^{(0)}$, we may assume without loss of generality that $\mathscr{G}^{(0)}$ is compact, and hence $K\mathscr{G}$ is unital, by Propositions 3.2 and 3.4.
Note that if $T$ is any inverse subsemigroup of $S$, then $\bigcup T$ is an open subgroupoid of $\mathscr{G}$. Also, we have that $\mathscr{G}=\bigcup S$. Since $S$ is a directed union of its finitely generated inverse subsemigroups, it suffices by Proposition 3.4 to prove that $\bigcup T$ is uniformly bounded for every finitely generated inverse subsemigroup $T$ of $S$. So let $X\subseteq S$ be a finite subset closed under inversion. We will show that $\mathscr{H}=\bigcup\langle X\rangle$ is uniformly bounded.
Say $X=\{U_{1},\ldots,U_{n}\}$ and let $a=\frac{1}{2n}(1_{U_{1}}+\cdots+1_{U_{n}})\in K\mathscr{G}$. Then since $K\mathscr{G}$ is regular, we can find $b\in K\mathscr{G}$ such that $(1-a)b(1-a)=(1-a)$. We can write $b=c_{1}1_{V_{1}}+\cdots+c_{k}1_{V_{k}}$ with $V_{1},\ldots,V_{k}$ compact open bisections. Then $L=\mathbb{Q}(c_{1},\ldots,c_{k})$ is a finitely generated extension of $\mathbb{Q}$ such that $a,b\in L\mathscr{G}$. Since $L$ is countable, we can embed it over $\mathbb{Q}$ into $\mathbb{C}$ and hence $a,b\in\mathbb{C}\mathscr{G}$. Note that by Proposition 3.6, we have that $\|a\|_{r}\leq|a|\leq\frac{1}{2}$, and so $\sum_{m=0}^{\infty}a^{m}$ converges to $(1-a)^{-1}$ in $C^{*}_{r}(\mathscr{G})$. But in $\mathbb{C}\mathscr{G}\subseteq C^{*}_{r}(\mathscr{G})$, we have $(1-a)b(1-a)=(1-a)$, and so we deduce that $b=(1-a)^{-1}$ by cancelling $1-a$ on both sides in $C^{*}_{r}(\mathscr{G})$. Thus $b=\sum_{m=0}^{\infty}a^{m}$ in $C^{*}_{r}(\mathscr{G})$, that is, $\lim_{N\to\infty}\||b-\sum_{m=0}^{N}a^{m}\|_{r}=0$. But then since $\|f\|_{\infty}\leq\|f\|_{r}$, for $f\in\mathbb{C}\mathscr{G}$ by Proposition 3.6, we have $\lim_{N\to\infty}\||b-\sum_{m=0}^{N}a^{m}\|_{\infty}=0$. From the definition of $a$, we see that
$a^{m}=\frac{1}{(2n)^{m}}\sum_{U_{i_{1}},\ldots,U_{i_{m}}\in X}1_{U_{i_{1}}\cdots U_{i_{m}}}.$
Now if $\gamma\in\mathscr{H}=\bigcup\langle X\rangle$, then $\gamma\in U_{i_{1}}\cdots U_{i_{r}}$ for some $U_{i_{1}},\ldots,U_{i_{r}}\in X$ (not necessarily distinct). But then $\sum_{m=0}^{N}a^{m}(\gamma)\geq a^{r}(\gamma)\geq\frac{1}{(2n)^{r}}$ for any $N\geq r$. It follows that $b(\gamma)\neq 0$ for all $\gamma\in\mathscr{H}$. Thus $\mathscr{H}$ is contained in the quasi-compact subspace $V_{1}\cup\cdots\cup V_{k}$ of $\mathscr{G}$. It follows from Lemma 3.3 that $\mathscr{H}$ is uniformly bounded, completing the proof. ∎
Remark 3.8.
Note that Theorem 3.7 gives us some additional information about the structure of $K\mathscr{G}$ when it is regular in characteristic 00. Namely, $K\mathscr{G}$ is a directed union of finite dimensional unital semisimple subalgebras. In particular, if $K$ is algebraically closed, then $K\mathscr{G}$ is ultramatricial (i.e., a direct limit of finite direct products of matrix algebras over $K$). Indeed, if $\Gamma_{c}(\mathscr{G})$ is locally finite, then since $KT$ is a finite dimensional semisimple $K$-algebra for any finite inverse subsemigroup of $T$ [32, Chapter 9], we see that $K\Gamma_{c}(\mathscr{G})$ is of the desired form. Since a quotient algebra of such a directed union is again such a directed union, it follows that $K\mathscr{G}$ is of this form.
In an attempt to understand the positive characteristic case, we will give sufficient conditions for the algebra of an approximately quasi-compact ample groupoid to be regular.
Proposition 3.9.
Let $\mathscr{G}$ be a quasi-compact ample group bundle and $K$ a field whose characteristic does not divide the order of any isotropy group element. Then $K\mathscr{G}$ is regular.
Proof.

### v3

@@ 2.4. Regular rings
A ring $A$ is von Neumann regular, or regular for short, if for all $a\in A$, there exists $b\in A$ with $aba=a$. We do not require rings to be unital, however the coefficient rings of our group, semigroup and groupoid algebras shall always be assumed unital. The following facts are quite standard; see [19, Chapter 1].
@@
    An arbitrary direct product of regular rings is regular.
@@
    If $A$ is a ring and $I$ is an ideal of $A$, then $A$ is regular if and only if $I$ and $A/I$ are regular.
@@
    A ring $A$ is regular if and only if it has local units and each unitary $A$-module is flat.
A semigroup is locally finite if all its finitely generated subsemigroups are finite, or, equivalently, it is a direct limit of finite semigroups. Note that any subsemigroup of a locally finite semigroup is locally finite. The following is a special case of a classical theorem of Birkhoff in universal algebra, cf. [8, Theorem 10.16]. We include a proof for the sake of completeness.
@@
Let $G$ be a group and $R$ a unital ring. Then the group ring $RG$ is regular if and only if $R$ is regular, $G$ is locally finite and the order of each finite subgroup of $G$ is invertible in $R$.
Our main result extends Connell’s theorem to ample groupoids.
We shall need the following result going back to Weissglass [34]. We sketch a proof for completeness.
Proposition 2.4 (Weissglass).
Let $R$ be a regular unital ring and let $S$ a locally finite inverse semigroup such that if a prime $p$ is the order of a cyclic subgroup of $S$, then $p$ is invertible in $R$. Then $RS$ is regular.
Proof.
Since the class of regular rings is closed under direct limits by Proposition 2.1, we may assume that $S$ is a finite inverse semigroup. Then it is well known that $RS$ is isomorphic to a finite direct product of matrix algebras over the group rings of its maximal subgroups (cf. [32, Theorem 4.6] where the assumption that $R$ is commutative is superfluous). Our assumptions imply that the order of any maximal subgroup $G$ of $S$ is invertible in $R$ by Cauchy’s theorem, and so $RG$ is regular by Theorem 2.3. Thus $RS$ is regular by Proposition 2.1. ∎
3. von Neumann regularity of Steinberg algebras
We begin by establishing the necessity of the conditions in Theorem A.
@@
Suppose that $R\mathscr{G}$ is regular with $\mathscr{G}$ ample and $R$ a unital ring. Then $R$ is regular.
@@
Let $r\in R$ and let $U$ be any nonempty compact open subset of $\mathscr{G}^{(0)}$. Let $f=r1_{U}$. Then there exists $g\in R\mathscr{G}$ with $fgf=f$. Let $x\in U$. Then $r=f(x)=fgf(x)=rg(x)r$, as $x=\alpha\beta\gamma$ with $\alpha,\gamma\in U$ implies $\alpha=x=\gamma$, and hence $\beta=x$. We conclude that $R$ is regular. ∎
Let us say that an ample groupoid $\mathscr{G}$ is uniformly bounded if there is an integer $M\geq 1$ with $|\mathop{\boldsymbol{r}}\nolimits^{-1}(x)|\leq M$ for all $x\in\mathscr{G}^{(0)}$ or, equivalently, $|\mathop{\boldsymbol{s}}^{-1}(x)|\leq M$ for all $x\in\mathscr{G}^{(0)}$. Note that any open or closed subgroupoid of a uniformly bounded ample groupoid is uniformly bounded, and so, in particular, a uniformly bounded groupoid has finite isotropy groups. Let us say that $\mathscr{G}$ is approximately uniformly bounded if it is a directed union of uniformly bounded open subgroupoids. An open or closed subgroupoid of an approximately uniformly bounded ample groupoid is again approximately uniformly bounded, and hence the isotropy groups of such a groupoid are locally finite. Let us say that $\mathscr{G}$ is approximately quasi-compact if $\mathscr{G}$ is a directed union of quasi-compact open subgroupoids. It is not difficult to verify that a principal second countable Hausdorff ample groupoid is approximately quasi-compact if and only if it is an approximately finite groupoid in the sense of [28, Section 11.5] (i.e., an approximately finite étale equivalence relation in the sense of [18]). Note that if $\mathscr{G}$ is a quasi-compact ample groupoid, then $\mathscr{G}^{(0)}$ is compact being the image of $\mathscr{G}$ under $\mathop{\boldsymbol{r}}\nolimits$.
@@
Let $C$ be a quasi-compact subspace of $\mathscr{G}$ containing $\mathscr{H}$. We may cover $C$ by finitely many compact open bisections $U_{1}\cup\cdots\cup U_{n}$ of $\mathscr{G}$. Then we have $|\mathop{\boldsymbol{r}}\nolimits^{-1}(x)\cap\mathscr{H}|\leq n$ for all $x\in\mathscr{H}^{(0)}$ as $|\mathop{\boldsymbol{r}}\nolimits^{-1}(x)\cap U_{i}|\leq 1$ for $i=1,\ldots,n$ and $\mathscr{H}\subseteq U_{1}\cup\cdots\cup U_{n}$. ∎
@@
Suppose that (2) holds. If $\mathscr{G}$ is the directed union of uniformly bounded open subgroupoids $\mathscr{G}_{\alpha}$ with $\alpha\in D$, then $\Gamma_{c}(\mathscr{G})$ is the directed union of the $\Gamma_{c}(\mathscr{G}_{\alpha})$ since any compact set belongs to some $\mathscr{G}_{\alpha}$. Hence any finitely generated subsemigroup of $\Gamma_{c}(\mathscr{G})$ is contained in some $\Gamma_{c}(\mathscr{G}_{\alpha})$. Thus we may assume without loss of generality that $\mathscr{G}$ is uniformly bounded. Suppose that $|\mathop{\boldsymbol{r}}\nolimits(x)|\leq n$ for all $x\in\mathscr{G}^{(0)}$. Then, for each $x\in\mathscr{G}^{(0)}$, we have that $|\mathscr{G}^{x}_{x}|\cdot|\mathcal{O}_{x}|=|\mathop{\boldsymbol{r}}\nolimits^{-1}(x)|\leq n$. It follows that $|\mathscr{G}|_{\mathcal{O}_{x}}|=|\mathcal{O}_{x}|^{2}\cdot|\mathscr{G}_{x}^{x}|\leq n^{2}$, and hence $\mathscr{G}|_{\mathcal{O}_{x}}$ is a finite closed subgroupoid, and $\Gamma_{c}(\mathscr{G}|_{\mathcal{O}_{x}})$ is a finite inverse semigroup of cardinality at most $2^{n^{2}}$.
@@
Now we prove conditions (2) and (3) of Theorem A are necessary.
@@
Let $R$ be a unital ring and $\mathscr{G}$ an ample groupoid such that $R\mathscr{G}$ is von Neumann regular. Then $\mathscr{G}$ is approximately quasicompact and the order of any finite subgroup of an isotropy group of $\mathscr{G}$ is invertible in $R$.
@@
First note that if $U$ is a compact open subset of $\mathscr{G}^{(0)}$, then $\mathscr{G}|_{U}$ is an open subgroupoid with compact unit space and $R\mathscr{G}|_{U}=1_{U}R\mathscr{G}1_{U}$. Hence if $R\mathscr{G}$ is regular, then so is $R\mathscr{G}|_{U}$ by Proposition 2.1. Since $\mathscr{G}$ is the directed union of open subgroupoids of the form $\mathscr{G}|_{U}$, we may assume without loss of generality that $\mathscr{G}^{(0)}$ is compact, that is, $R\mathscr{G}$ has the identity element $1_{\mathscr{G}^{(0)}}$.
By Proposition 3.4, it suffices to show that $\mathscr{G}$ is approximately uniformly bounded, and that the order of any finite subgroup of an isotropy group is invertible in $R$. If $S\leq\Gamma_{c}(\mathscr{G})$ is an inverse submonoid, then $\bigcup S$ is an open subgroupoid of $\mathscr{G}$ containing $\mathscr{G}^{(0)}$. Moreover, since $\Gamma_{c}(\mathscr{G})$ is the directed union of its finitely generated inverse submonoids and $\mathscr{G}=\bigcup\Gamma_{c}(\mathscr{G})$, it suffices to show that if $S\leq\Gamma_{c}(\mathscr{G})$ is a finitely generated inverse submonoid, then $\mathscr{H}=\bigcup S$ is uniformly bounded and the order of any isotropy group of $\mathscr{H}$ is invertible in $R$. We proceed by adapting the idea of Connell for the case of groups [12].
Let $I_{S}$ be the left ideal of $R\mathscr{G}$ generated by the elements $1_{U}-1_{UU^{-1}}$ with $U\in S$. Let $X$ be a finite generating set of $S$ as a semigroup which is closed under taking inverses. We prove that $I_{S}$ is generated as a left ideal by the finite set $Y$ of elements of the form $1_{V}-1_{VV^{-1}}$ with $V\in X$. We proceed by induction on word length, i.e., the minimum length of an expression of an element $U$ of $S$ as a product of elements of $X$, to show that $1_{U}-1_{UU^{-1}}\in K\mathscr{G}\cdot Y$, with the case of length $1$ being trivial. If $U=WV$ with $V\in X$ and $W\in S$ of length one shorter than $U$, then $W^{-1}$ also has length one shorter than $U$ (since $X$ is closed under inversion), and so
$1_{U}-1_{UU^{-1}}=1_{W}(1_{V}-1_{VV^{-1}})-1_{WVV^{-1}}(1_{W^{-1}}-1_{W^{-1}W})\in K\mathscr{G}\cdot Y$
by induction, as required.
Since $R\mathscr{G}$ is a regular ring with identity and $I_{S}$ is a finitely generated left ideal, there is an idempotent $e\in R\mathscr{G}$ with $I_{S}=R\mathscr{G}e$ by [19, Theorem 1.1]. Let $f=1_{\mathscr{G}^{(0)}}-e$. Then $f$ is an idempotent and $I_{S}f=R\mathscr{G}e(1-e)=0$, that is, $1_{U}f=1_{UU^{-1}}f$ for all $U\in S$. If $f=\sum_{i=1}^{r}c_{i}1_{U_{i}}$ with $U_{i}$ compact open bisections, then $f$ can be nonzero on at most $r$ elements from the same $\mathop{\boldsymbol{s}}$-fiber, a fact that we shall soon use.
Let $\gamma\in\mathscr{H}$ and $\beta\in\mathscr{G}$ with $\mathop{\boldsymbol{s}}(\gamma)=\mathop{\boldsymbol{r}}\nolimits(\beta)$. We claim that $f(\gamma\beta)=f(\beta)$. Indeed, let $U\in S$ with $\gamma\in U$. Then $(1_{U}f)(\gamma\beta)=f(\beta)$ and $(1_{UU^{-1}})f(\gamma\beta)=f(\gamma\beta)$. We conclude that $f(\gamma\beta)=f(\beta)$ for all $\gamma\in\mathscr{H}\cap\mathop{\boldsymbol{s}}^{-1}(\mathop{\boldsymbol{r}}\nolimits(\beta))$.
Let $\mathop{\boldsymbol{r}}\nolimits_{*}\colon R\mathscr{G}\to R\mathscr{G}^{(0)}$ be the left $R$-module homomorphism induced by the local homeomorphism $\mathop{\boldsymbol{r}}\nolimits$. So
$\mathop{\boldsymbol{r}}\nolimits_{*}(f)(x)=\sum_{\mathop{\boldsymbol{r}}\nolimits(\gamma)=x}f(\gamma).$
Note that if $U\in\Gamma_{c}(\mathscr{G})$, then $\mathop{\boldsymbol{r}}\nolimits_{*}(1_{U})=1_{UU^{-1}}$. It is well known and easy to check that $\mathop{\boldsymbol{r}}\nolimits_{*}$ is a left $R\mathscr{G}$-module homomorphism where $R\mathscr{G}$ acts on $R\mathscr{G}^{(0)}$ via the rule
$f_{1}\cdot g_{1}(x)=\sum_{\mathop{\boldsymbol{r}}\nolimits(\gamma)=x}f_{1}(\gamma)g_{1}(\mathop{\boldsymbol{s}}(\gamma))$
for $f_{1}\in R\mathscr{G}$ and $g_{1}\in R\mathscr{G}^{(0)}$. Notice that if $U\in\Gamma_{c}(\mathscr{G})$ and $V\subseteq\mathscr{G}^{(0)}$ is compact open, then $1_{U}\cdot 1_{V}=1_{UVU^{-1}}$.
If $U\in S$, then $\mathop{\boldsymbol{r}}\nolimits_{*}(1_{U}-1_{UU^{-1}})=0$, and so $I_{S}\subseteq\ker\mathop{\boldsymbol{r}}\nolimits_{*}$. Therefore, $\mathop{\boldsymbol{r}}\nolimits_{*}(f)=\mathop{\boldsymbol{r}}\nolimits_{*}(1_{\mathscr{G}^{(0)}}-e)=\mathop{\boldsymbol{r}}\nolimits_{*}(1_{\mathscr{G}^{(0)}})=1_{\mathscr{G}^{(0)}}$, as $e\in I_{S}$. Let $x\in\mathscr{G}^{(0)}$. Then
$1=\mathop{\boldsymbol{r}}\nolimits_{*}(f)(x)=\sum_{\mathop{\boldsymbol{r}}\nolimits(\beta)=x}f(\beta)$ (3.1)
and so we can find $\beta\in\mathscr{G}$ with $\mathop{\boldsymbol{r}}\nolimits(\beta)=x$ and $f(\beta)\neq 0$. Then, for every $\gamma\in\mathscr{H}\cap\mathop{\boldsymbol{s}}^{-1}(x)$, we have $f(\gamma\beta)=f(\beta)\neq 0$. Since $f$ is nonzero on at most $r$ elements with source $\mathop{\boldsymbol{s}}(\beta)$, we deduce that $|\mathop{\boldsymbol{s}}^{-1}(x)\cap\mathscr{H}|<r+1$. Finally, the group $\mathscr{H}_{x}^{x}$ acts freely on the left of $\mathop{\boldsymbol{r}}\nolimits^{-1}(x)$, and so if we choose a transversal $T$ for this action and use $f(\gamma\beta)=f(\beta)$ for $\gamma\in\mathscr{H}_{x}^{x}$, we obtain from (3.1)
$1=|\mathscr{H}_{x}^{x}|\cdot\sum_{t\in T}f(t),$
and so $|\mathscr{H}_{x}^{x}|$ is invertible in $R$. This completes the proof. ∎
We now prove the sufficiency of the conditions in Theorem A.
Proposition 3.7.
Let $\mathscr{G}$ be a quasi-compact ample group bundle and $R$ a regular unital ring such that the order of any cyclic subgroup of an isotropy group of $\mathscr{G}$ is invertible in $R$. Then $R\mathscr{G}$ is regular.
@@
We know that $S=\Gamma_{c}(\mathscr{G})$ is locally finite by Proposition 3.4. We show that $RS$ is regular. It will then follow that its quotient $R\mathscr{G}$ is regular by Proposition 2.1. By Proposition 2.4 it suffices to show that if a prime $p$ is the order of a cyclic subgroup of $S$, then $p$ is invertible in $R$.